OpenStudy (anonymous):

Which of the following is a solution of y > |x| - 6? (-5, 1) (-1, -5) (5, -1)

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

\[y+6>\left| x \right|or \left| x \right|<y+6\] -(y-6)<x<y+6

OpenStudy (anonymous):

\[y+6>\left| x \right|or \left| x \right|<y+6\] -(y-6)<x<y+6

OpenStudy (anonymous):

What? @_@

OpenStudy (anonymous):

for this I would plug in each solution into the equation and see which one holds true, so like this: y > abs(x) -6 check with (5,-1) then -1>abs(5)-6 -1>5-6 -1>-1 [False] do this for the the set (-5, 1) and (-1, -5) and see which one is True ^_^

OpenStudy (anonymous):

i think its (-5,1)

OpenStudy (anonymous):

is it @DemolisionWolf

OpenStudy (anonymous):

correct ^_^

OpenStudy (anonymous):

yay lol